Clarita borrowed money at a 6% interest rate. If sheh pays the bank a total of $8000 at the end of ten years, how much did she originally borrow?
How do I utilize the I=prt formula to figure out the original borrowing amount?
Thanks!

Answer by geetha_rama(94)

Amount = Principle + Interest -> formula 1
Interest = (P*r*t)/100 -> formula 2
Let P be the Principle that Clarita borrowed from the bank
I be the interest
r = rate of interest = 6%
t = 10 years
Amount = $8000
Using the formula 1,
8000 = P + I
Substituting I from the formula 2,
8000 = P + (Prt)/100]
8000 = P*(1+ (rt)/100)
8000 = P*(1+ (6* 10)/100)
8000 = P*(1+3/5)
8000= P*8/5
=> P = (8000*5)/8
=> P = 5000
Hence Clarita borrowed $5000 from the bank